OpenStudy (anonymous):

. Solve by simulating the problem. You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly? Type your answer below using complete sentences

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

This is a binomial distribution.

OpenStudy (anonymous):

@shamil98 This is good practice for you as well

OpenStudy (shamil98):

n = 5 k = 3

OpenStudy (anonymous):

And p?

OpenStudy (shamil98):

p = 3/5 ?

OpenStudy (anonymous):

Nope. A trial here is any multiple choice question.

OpenStudy (anonymous):

If you randomly guess 1 of the 4 questions, what is probability of said trial winning?

OpenStudy (shamil98):

1/4?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Do you think you can explain it to Jasmine?

OpenStudy (anonymous):

If not I'll have to explain it a bit later.

OpenStudy (shamil98):

Alright so, jasmine, we plug in n k and p into the following formula f(k) = n! / k!(n-k)! p^k (1-p)^n-k

OpenStudy (shamil98):

We have identified n k and p. So, what would the equation look like?

OpenStudy (anonymous):

f(3)=5 / 3(5-3) 1/4 ^3 (1-1/4) ^5-3

OpenStudy (shamil98):

Do you know what factorials are? 5! for example?

OpenStudy (shamil98):

\[f(3) =\frac{ 5! }{ 3!(5-3)!} (1/4)^3 (1-1/4)^2\]

OpenStudy (shamil98):

\[5! = 5 \times 4 \times 3 \times 2 \times 1 = 120\]

OpenStudy (shamil98):

A factorial is basically every consecutive number up to that number for example.. \[3! = 3 \times 2 \times 1 = 6\]

OpenStudy (shamil98):

\[f(3) = \frac{ 120 }{ 6 \times 2! } (1/4)^3 (1-1/4)^2\]

OpenStudy (dan815):